OpenStudy (anonymous):

“To pass by me, you must tell me how to convert standard form into the general, vertex form... I have a test on it next week.” Explain how to convert f(x) into the general, vertex form of the equation. Use complete sentences.

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

You need to complete the square.

OpenStudy (mathstudent55):

Standard form is: f(x) = ax^2 + bx + c Vertex form is f(x) = a(x - h)^2 + k